# What Topological and Geometric Structure Do Biological Foundation Models Learn? Evidence from 141 Hypotheses

### v1

@@ 1 Introduction
Foundation models for single-cell genomics have demonstrated impressive performance on downstream tasks, from cell-type annotation to gene perturbation prediction (Cui et al., 2024; Theodoris et al., 2023). But an unsettling question lingers beneath these successes: what do these models actually understand about biology? Are their internal representations organized in biologically meaningful ways, or are they opaque statistical summaries that happen to correlate with biological outcomes?
The stakes are high. If these models encode meaningful biological structure internally—regulatory relationships, protein interactions, pathway organization—then their representations could be mined for biological discovery. If not, claims about “biological knowledge” in foundation models rest on fragile ground.
Mechanistic interpretability research in language models has shown that neural representations can be surprisingly structured, encoding linear concept representations (Park et al., 2023), feature superposition (Elhage et al., 2022), and depth-dependent processing stages (Nanda et al., 2023). In the biological domain, recent work has examined this question through two complementary lenses. First, analysis of attention patterns in scGPT and Geneformer revealed that attention encodes protein-protein interactions in early layers and transcriptional regulation in later layers—but provides no incremental value for perturbation prediction beyond trivial gene-level baselines (Kendiukhov, 2025). Second, spectral analysis of residual-stream geometry showed that scGPT organizes genes along interpretable axes corresponding to subcellular localization, interaction networks, and regulatory identity (Kendiukhov, 2026).
These studies established that biological transformers encode linear geometric structure. But a fundamental question remains: does the embedding space contain deeper nonlinear geometric and topological structure—loops, manifold curvature, community boundaries, directional asymmetries—that carries biological meaning beyond what linear analysis captures? And critically, how much of any apparent structure survives rigorous null controls?
@@ 2.2 Models and Data
We analyzed scGPT (Cui et al., 2024) (12 transformer layers) by extracting residual-stream gene embeddings: for each tissue context, single-cell expression profiles from the Tabula Sapiens atlas (The Tabula Sapiens Consortium, 2022) were fed through the pretrained model, and the hidden-state vectors at each transformer layer were averaged across cells to produce a single embedding vector per gene per layer. Three tissue domains were used: lung, immune, and external-lung (a held-out lung dataset not used for any parameter tuning), each with three independent random seeds controlling gene sampling and train/test partitioning. For cross-model comparisons, we extracted analogous gene embeddings from Geneformer V2-316M (Theodoris et al., 2023) (18 layers, 18 heads) on the same tissue domains, restricting to genes present in both models’ vocabularies (typically 320 shared genes per domain). Typical experiments sampled 160–350 genes per test, with PCA dimensionality reduction (10–20 components) applied to the embedding vectors for computational tractability and to mitigate the curse of dimensionality in distance computations.
Regulatory ground truth was drawn from DoRothEA (Garcia-Alonso et al., 2019) (confidence-tiered TF–target annotations), TRRUST (Han et al., 2018) (signed regulatory edges indicating activation or repression), STRING (Szklarczyk et al., 2019) (protein–protein interaction confidence scores), and Gene Ontology (Ashburner et al., 2000) (functional co-membership annotations).
@@ 3.1 Two Independently Trained Models Converge on the Same Geometric Organization
We tested this with progressively refined alignment methods. Canonical correlation analysis with whitening (H24)—which projects both models’ PCA-reduced gene embeddings into a shared subspace that maximizes inter-model correlation, then evaluates whether geometric relationships are preserved—revealed remarkable geometric consistency between matched scGPT and Geneformer gene sets: mean canonical correlation of 0.80, pairwise-distance Spearman correlation of 0.75, and gene-level top-1 retrieval accuracy of 72% (combined Fisher across all three tissue domains). Simpler Procrustes alignment (H20) confirmed the pattern: 40% top-1 retrieval accuracy, significant in all domains (). Even the ranking of which geometric features are most informative was conserved across models (H17): centered cosine was the top-ranked feature in all three domains for both models ().
@@ 3.2 Gene Embedding Neighborhoods Contain Non-Trivial Topology
We next asked whether scGPT’s gene representations contain topological structure—specifically, whether the embedding neighborhoods form “loops” beyond what would be expected from random geometric arrangements. Persistent homology detects such loops by constructing a filtration of simplicial complexes over increasing distance thresholds and identifying one-dimensional topological features (“H1 classes”) that persist across a range of thresholds; the total persistence (sum of birth-to-death lifetimes across all H1 classes) quantifies how much loop-like structure is present.
The answer is clearly yes, under appropriate controls. Computing H1 persistence (via the Ripser algorithm) on PCA-projected embedding neighborhoods (350 genes, PCA to 20 dimensions) and comparing against feature-shuffle nulls (20 replicates per condition), we found that scGPT representations in the lung domain have significantly more topological structure than shuffled embeddings in 11 of 12 transformer layers ( in 9/12; top layer L0: , ). This signal replicated across all three tissue domains: immune (12/12 layers significant, mean H1 persistence delta lifetime units) and external-lung (12/12 layers significant, mean delta ).
@@ 3.3 The Embedding Manifold Encodes a Distance Hierarchy for Regulation
A clear hierarchy emerged. Euclidean distance (straight-line in embedding space) was outperformed by geodesic distance (shortest path along the NN manifold, with adaptively chosen from 10–35 to ensure graph connectivity; AUROC , 7/12 layers significant in both gene-pool splits). Geodesic distance was in turn outperformed by diffusion distance: the distance between rows of a diffusion-map embedding derived from a random walk on the NN graph, with the diffusion time swept over and selected by best performance (AUROC , significant in all three domains; Fisher -values from to ). Convexity-deficit and detour-based graph-geometry metrics provided additional evidence for the distance hierarchy (H32: AUROC , Fisher-significant in 4/6 domain-splits). The strongest single geometry metric was the multiscale triangle-defect spectrum (H69/H70). For each gene triplet in a NN neighborhood, the triangle defect measures how much the pairwise distances deviate from the triangle inequality—quantifying local departure from flat Euclidean geometry. Aggregating these defects at multiple neighborhood scales () produces a per-edge feature vector that captures multiscale curvature around each gene pair. This metric achieved AUROC with positive null-gap in all domain-split groups under hard-null calibration. This metric became the foundation for all subsequent work in the campaign: nearly every hypothesis from H70 onward was evaluated as an incremental improvement over this triangle-defect baseline.
@@ 3.4 The Strongest Finding: Regulatory Motifs Align with Geometric Community Structure
The single most robust finding across the entire 141-hypothesis screen emerged from combining two sources of information: the geometric community structure of the embedding space (Louvain communities detected on the NN graph of gene embeddings) and the signed regulatory motif annotations from the TRRUST database (which transcription factors activate or repress which targets).
@@ 3.5 Combining Geometric Features Through Stability Selection
No single geometric measure captured all the regulatory information present in the embeddings. This motivated a multivariate approach: for each gene pair, we computed a vector of geometric edge features (geodesic distance, triangle defects, community co-membership, bifiltration cycle-rank, directed topology features) and combined them via stability selection—a procedure that fits randomized LASSO models on many bootstrap subsamples and retains only features selected in a high fraction of runs, yielding a sparse, reproducible feature set that is then used in a cross-validated logistic classifier.
@@ 3.7 What the Models Do Not Encode: 70+ Negative Results
| Forman curvature enrichment (H23) | High-curvature edges are less likely to be regulatory (AUROC 0.34–0.39, below chance)—opposite to the hypothesis |
@@
Across 19 distinct methods—Gromov–Wasserstein transport, seeded and unseeded optimal transport, topology-signature distillation, cycle-consistent mapping, codebook transport, chart/sheaf alignment, and more—the mean top-1 retrieval rate for gene-level correspondence was below 1%. Only perturbation-response alignment (H108) showed partial promise, but it failed to replicate across seeds (H109). This is not a failure of methodology; it is an intrinsic property of these models. They learn similar relationships between genes (distances, neighborhoods, clusters) but embed individual genes at different coordinates within that shared relational structure.
@@ 4.4 Limitations
Our screening covered three tissue domains from a single species; extension to additional tissues, cell types, and species is needed to assess generality. The regulatory ground truth databases (DoRothEA, TRRUST, STRING) are incomplete and biased toward well-studied gene families, so some “negative” results may reflect annotation limitations. The autonomous loop used a greedy exploration strategy; more systematic approaches (e.g., Bayesian optimization) might discover additional positive signals. Finally, whether the geometric structure identified here translates to practical value for downstream biological tasks—perturbation prediction, drug target prioritization, pathway discovery—remains to be tested.
@@ 5 Conclusion
All experimental scripts, iteration reports, raw CSV results, and the autonomous loop orchestration code are available in the project repository at https://github.com/Biodyn-AI/topology-biomechinterp2.

### v2

@@ 1 Introduction
Foundation models for single-cell genomics have demonstrated impressive performance on downstream tasks, from cell-type annotation to gene perturbation prediction (Cui et al., 2024; Theodoris et al., 2023; Yang et al., 2022; Hao et al., 2024). But an unsettling question lingers beneath these successes: what do these models actually understand about biology? Are their internal representations organized in biologically meaningful ways, or are they opaque statistical summaries that happen to correlate with biological outcomes?
The stakes are high. If these models encode meaningful biological structure internally—regulatory relationships, protein interactions, pathway organization—then their representations could be mined for biological discovery (Lopez et al., 2018; Luecken and Theis, 2019). If not, claims about “biological knowledge” in foundation models rest on fragile ground.
Mechanistic interpretability research in language models has shown that neural representations can be surprisingly structured, encoding linear concept representations (Park et al., 2023), feature superposition (Elhage et al., 2022), depth-dependent processing stages (Nanda et al., 2023), and recoverable computational circuits (Conmy et al., 2023; Olah et al., 2020). In the biological domain, recent work has examined this question through two complementary lenses. First, analysis of attention patterns in scGPT and Geneformer revealed that attention encodes protein-protein interactions in early layers and transcriptional regulation in later layers—but provides no incremental value for perturbation prediction beyond trivial gene-level baselines (Kendiukhov, 2025). Second, spectral analysis of residual-stream geometry showed that scGPT organizes genes along interpretable axes corresponding to subcellular localization, interaction networks, and regulatory identity (Kendiukhov, 2026).
These studies established that biological transformers encode linear geometric structure. But a fundamental question remains: does the embedding space contain deeper nonlinear geometric and topological structure—loops, manifold curvature, community boundaries, directional asymmetries—that carries biological meaning beyond what linear analysis captures (Carlsson, 2009; Rabadan and Blumberg, 2019)? And critically, how much of any apparent structure survives rigorous null controls?
@@ 2.2 Models and Data
We analyzed scGPT (Cui et al., 2024) (12 transformer layers (Vaswani et al., 2017)) by extracting residual-stream gene embeddings: for each tissue context, single-cell expression profiles from the Tabula Sapiens atlas (The Tabula Sapiens Consortium, 2022) were fed through the pretrained model, and the hidden-state vectors at each transformer layer were averaged across cells to produce a single embedding vector per gene per layer. Three tissue domains were used: lung, immune, and external-lung (a held-out lung dataset not used for any parameter tuning), each with three independent random seeds controlling gene sampling and train/test partitioning. For cross-model comparisons, we extracted analogous gene embeddings from Geneformer V2-316M (Theodoris et al., 2023) (18 layers, 18 heads) on the same tissue domains, restricting to genes present in both models’ vocabularies (typically 320 shared genes per domain). Typical experiments sampled 160–350 genes per test, with PCA dimensionality reduction (10–20 components) applied to the embedding vectors for computational tractability and to mitigate the curse of dimensionality in distance computations.
Regulatory ground truth was drawn from DoRothEA (Garcia-Alonso et al., 2019) (confidence-tiered TF–target annotations), TRRUST (Han et al., 2018) (signed regulatory edges indicating activation or repression), STRING (Szklarczyk et al., 2019) (protein–protein interaction confidence scores), and Gene Ontology (Ashburner et al., 2000; Gene Ontology Consortium, 2021) (functional co-membership annotations). Gene regulatory network inference from single-cell data remains a major benchmark challenge (Pratapa et al., 2020; Dibaeinia and Sinha, 2020), and the evaluation design adopted here reflects lessons from the benchmarking literature on the sensitivity of method rankings to protocol choices (Saelens et al., 2019).
@@ 3.1 Two Independently Trained Models Converge on the Same Geometric Organization
We tested this with progressively refined alignment methods. Canonical correlation analysis (Hardoon et al., 2004) with whitening (H24)—which projects both models’ PCA-reduced gene embeddings into a shared subspace that maximizes inter-model correlation, then evaluates whether geometric relationships are preserved—revealed remarkable geometric consistency between matched scGPT and Geneformer gene sets: mean canonical correlation of 0.80, pairwise-distance Spearman correlation of 0.75, and gene-level top-1 retrieval accuracy of 72% (combined Fisher across all three tissue domains). Simpler Procrustes alignment (H20) confirmed the pattern: 40% top-1 retrieval accuracy, significant in all domains (). Even the ranking of which geometric features are most informative was conserved across models (H17): centered cosine was the top-ranked feature in all three domains for both models ().
@@ 3.2 Gene Embedding Neighborhoods Contain Non-Trivial Topology
We next asked whether scGPT’s gene representations contain topological structure—specifically, whether the embedding neighborhoods form “loops” beyond what would be expected from random geometric arrangements. Persistent homology (Edelsbrunner and Harer, 2008; Zomorodian and Carlsson, 2005) detects such loops by constructing a filtration of simplicial complexes over increasing distance thresholds and identifying one-dimensional topological features (“H1 classes”) that persist across a range of thresholds; the total persistence (sum of birth-to-death lifetimes across all H1 classes) quantifies how much loop-like structure is present.
The answer is clearly yes, under appropriate controls. Computing H1 persistence (via the Ripser algorithm (Bauer, 2021)) on PCA-projected embedding neighborhoods (350 genes, PCA to 20 dimensions) and comparing against feature-shuffle nulls (20 replicates per condition), we found that scGPT representations in the lung domain have significantly more topological structure than shuffled embeddings in 11 of 12 transformer layers ( in 9/12; top layer L0: , ). This signal replicated across all three tissue domains: immune (12/12 layers significant, mean H1 persistence delta lifetime units) and external-lung (12/12 layers significant, mean delta ).
@@ 3.3 The Embedding Manifold Encodes a Distance Hierarchy for Regulation
A clear hierarchy emerged. Euclidean distance (straight-line in embedding space) was outperformed by geodesic distance (shortest path along the NN manifold, with adaptively chosen from 10–35 to ensure graph connectivity; AUROC , 7/12 layers significant in both gene-pool splits). Geodesic distance was in turn outperformed by diffusion distance (Coifman and Lafon, 2006): the distance between rows of a diffusion-map embedding derived from a random walk on the NN graph, with the diffusion time swept over and selected by best performance (AUROC , significant in all three domains; Fisher -values from to ). Convexity-deficit and detour-based graph-geometry metrics provided additional evidence for the distance hierarchy (H32: AUROC , Fisher-significant in 4/6 domain-splits). The strongest single geometry metric was the multiscale triangle-defect spectrum (H69/H70). For each gene triplet in a NN neighborhood, the triangle defect measures how much the pairwise distances deviate from the triangle inequality—quantifying local departure from flat Euclidean geometry. Aggregating these defects at multiple neighborhood scales () produces a per-edge feature vector that captures multiscale curvature around each gene pair. This metric achieved AUROC with positive null-gap in all domain-split groups under hard-null calibration. This metric became the foundation for all subsequent work in the campaign: nearly every hypothesis from H70 onward was evaluated as an incremental improvement over this triangle-defect baseline.
@@ 3.4 The Strongest Finding: Regulatory Motifs Align with Geometric Community Structure
The single most robust finding across the entire 141-hypothesis screen emerged from combining two sources of information: the geometric community structure of the embedding space (Louvain communities (Blondel et al., 2008) detected on the NN graph of gene embeddings) and the signed regulatory motif annotations from the TRRUST database (which transcription factors activate or repress which targets).
@@ 3.5 Combining Geometric Features Through Stability Selection
No single geometric measure captured all the regulatory information present in the embeddings. This motivated a multivariate approach: for each gene pair, we computed a vector of geometric edge features (geodesic distance, triangle defects, community co-membership, bifiltration cycle-rank, directed topology features) and combined them via stability selection (Meinshausen and Bühlmann, 2010)—a procedure that fits randomized LASSO models on many bootstrap subsamples and retains only features selected in a high fraction of runs, yielding a sparse, reproducible feature set that is then used in a cross-validated logistic classifier.
@@ 3.7 What the Models Do Not Encode: 70+ Negative Results
| Forman curvature (Sreejith et al., 2016) enrichment (H23) | High-curvature edges are less likely to be regulatory (AUROC 0.34–0.39, below chance)—opposite to the hypothesis |
@@
Across 19 distinct methods—Gromov–Wasserstein transport (Peyré and Cuturi, 2019), seeded and unseeded optimal transport, topology-signature distillation, cycle-consistent mapping, codebook transport, chart/sheaf alignment, and more—the mean top-1 retrieval rate for gene-level correspondence was below 1%. Only perturbation-response alignment (H108) showed partial promise, but it failed to replicate across seeds (H109). This is not a failure of methodology; it is an intrinsic property of these models. They learn similar relationships between genes (distances, neighborhoods, clusters) but embed individual genes at different coordinates within that shared relational structure.
@@ 4.4 Limitations
Our screening covered three tissue domains from a single species; extension to additional tissues, cell types, and species is needed to assess generality. The regulatory ground truth databases (DoRothEA, TRRUST, STRING) are incomplete and biased toward well-studied gene families (Schaefer et al., 2012; Aibar et al., 2017), so some “negative” results may reflect annotation limitations. The autonomous loop used a greedy exploration strategy; more systematic approaches (e.g., Bayesian optimization) might discover additional positive signals. Finally, whether the geometric structure identified here translates to practical value for downstream biological tasks—perturbation prediction, drug target prioritization, pathway discovery—remains to be tested.
@@ 5 Conclusion
All experimental scripts, iteration reports, raw CSV results, and the autonomous loop orchestration code are available in the project repository at https://github.com/ihorkendiukhov/biomechinterp.